# Model Predictive Adversarial Imitation Learning for Planning from Observation

### v1

@@ 4 Experimental Results
Expectations. At a high level, MPAIL can be viewed as a model-based RL algorithm with a simultaneous expert-distribution matching objective. If provided a model, one would expect MPAIL to be more sample-efficient yet have worse asymptotic performance than current model-free methods like GAIL and AIRL [32].
Questions. Without a policy network, the conventional generator (i.e., policy) improvement step does not exist. In fact, the distinction between the discriminator and the policy blurs due to the direct discriminator costing in MPPI. While direct discriminator optimization is more akin to classical GAN networks [13], it is impractical to solve the optimal policy to completion under the current reward through planning as in IRL. This raises a central question regarding MPAIL’s training dynamics: Are approximated policies through value bootstrapping as in infinite-horizon MPPI sufficiently effective and stable to perform as adversarially generative policies?
In this section, we empirically validate that MPAIL indeed behaves as an AIL algorithm while providing the benefits of planning. We also aim to answer these follow-up questions in our experiments:
@@
    Is a learned AIL planner more robust than a learned AIL policy?
@@
    How can MPAIL help enable real-world planning capabilities from observation?
@@
4.1 Simulated Navigation Task
@@
For simulated evaluation, we design a navigation task with a 10-DoF vehicle. Reward is proportional to the negative squared distance to (10,10)(10,10)( 10 , 10 ). Initial poses ($t=0$) are within 1 m of (0,0)(0,0)( 0 , 0 ). The state is 12-dimensions: position, orientation, linear velocity, and angular velocity. Actions include target velocity and steering angle. MPAIL plans using an approximate prior model, the Kinematic Bicycle Model [33]. This approximate model is not tuned to the agent dynamics. For instance, slipping and suspension dynamics occur in simulation but are unmodeled.
@@
An expert is trained on this environment using PPO [34] and is used for expert data. At convergence, the optimal policy occasionally circles near the goal instead of stopping on it precisely (see Figure 9). While the total return between these two behaviors are nearly identical, we choose to use the circling demonstrations as expert data, because it is a more challenging behavior to imitate. This is corroborated by Orsini et al. [31], who stress that demonstrator suboptimality and multimodality is a critical component in algorithm evaluation towards practical AIL from human data.
@@
The expert’s circling behavior presents itself as a challenging “distractor mode”. In training, the policy may begin to only circle around the origin. If the AIL algorithm is not able to sufficiently explore, training collapses on this mode where the policy continuously circles the origin, unable to return to the expert distribution which requires the circling behavior to occur around the goal. For instance, we find that AIRL is unable to successfully learn both behaviors likely due to the instability introduced by its logit shift [31].
4.2 Out-of-Distribution Recovery Through Planning – Q1
@@
Figure 3: Comparison of policy-based and planning-based AIL in Out-of-Distribution (OOD) states. Agents trained on the navigation task (Section 4.1) are placed uniformly with random orientation between a 40 $\times$ 40 m box centered on (0,0)(0,0)( 0 , 0 ). The policy and planner are run for 100 timesteps in the environment. Data support of the expert exists mainly between (0,0)(0,0)( 0 , 0 ) and (10,10)(10,10)( 10 , 10 ) [^2] Note that the state space is actually 12-dimensional.. Quantitative evaluation of this experiment can be found in Figure 4.
@@
Figure 4: Quantitative OOD Comparison. Agent initial poses vary from In-distribution (ID) to OOD relative to the expert data and are plotted with their final reward after 100 timesteps.
@@
When deploying learning-based methods to the real-world, reliable performance in out-of-distribution (OOD) states are of critical importance. Especially in imitation learning when the expert data support can be extremely sparse, we show that a planning-based method like MPAIL can help improve generalization capabilities when OOD. In this experiment, we use the simulated navigation environment but expand the region of uniformly distributed initial positions and orientations from a 1$\times$1 square to a large 40$\times$40 m square around the origin (Footnote 2). The policy-based approach is represented by GAIL, as AIRL does not meaningfully converge in the navigation task (Section 4.1).
Planning-based AIL generalizes significantly better than policy-based AIL when outside both expert and training distributions. Figure 4 quantifies these findings through an energy-based OOD metric introduced by Liu et al. [35]. When viewing MPPI as an online policy optimizer, it should not be surprising that the MPAIL procedure provides benefits when OOD even if using an approximate model. However, this perspective requires the local optimization landscape as defined by the discriminator and value function to also generalize. Note that the planning horizon in this experiment is a maximum of 3 meters. These results suggest that model-based planning with value functions may indeed generalize better than policies for sparsely rewarded tasks like imitation learning. Meanwhile, there remain limits to the value function still. Stark OOD configurations around (15,15)(-15,-15)( - 15 , - 15 ) remains challenging for both value function and policy.
4.3 Real-Sim-Real Navigation from a Single Demonstration – Q2
@@
Real-world evaluation of AIL is currently challenging. RL sample inefficiency renders training in simulation more practical than in the real-world [36]. However, demonstrations must realistically be from the real-world. In addition, discriminative rewards may become entangled with the environment dynamics such that AIL will not produce a meaningful reward and thus policy [21].
Motivated by this gap, our hardware experiment involves: 1) collecting a single partially observable (position and body-centric velocity) demonstration from the real-world, 2) training the method through interactions from simulation, and 3) returning to the real-world to deploy the trained method. This experiment uses a small-scale RC car platform with an NVIDIA Jetson Orin NX [37]. The IRL-MPC baseline represents the approach of learning a reward and value using model-free AIL (GAIL), then using a hand-tuned MPPI to optimize these objectives on deployment.
We find that MPAIL is able to qualitatively reproduce the expert trajectory with an average Relative Cross-Track-Error (CTE [38]) of 0.17 m while traveling an average of 0.3 m/s slower. In addition, Figure 5 illustrates a key advantage of planning-based AIL. By granting access to the agent’s optimization landscape, MPAIL significantly improves on the interpretability of agents trained through ambiguous and complex human demonstration data when compared to black-box policies. Note the lower costing of on-track trajectories and final plan.
GAIL does not reliably converge to the expert even in training. During deployment, GAIL’s policy consistently veers off-path or collapses into driving in a circle. Various starting configurations were attempted without success. While literature on the evaluation of AIL methods in the real-world are sparse, we find that AIL policies can be extremely poor performing in the real-world, as corroborated by [19]. A more detailed discussion is provided in Section D.1.
@@
Table 1: Evaluation of Real Experiment. Relative Cross-Track Error (CTE) and speed are computed over the best five laps.
IRL-MPC is an elucidating middle-ground between MPAIL and GAIL; the learned reward and value are exactly the same as GAIL’s and thus differs mainly by the inclusion of a planner during deployment. IRL-MPC’s partial improvement over GAIL provides evidence that: (i) model-based planning can grant robustness to a model-free reward and, (ii) despite GAIL’s failure to produce meaningful behavior, the learned reward was still meaningfully discriminative and suggests a failure of the policy to perform as a competitive generator. With respect to MPAIL, IRL-MPC diverges by mainly learned reward and value. MPAIL’s improvements provides evidence that $\pi_{\text{MPPI}}$ induces a fundamentally different learning dynamic from offline policy optimization which supports MPAIL’s end-to-end formulation. In contrast to policy-based AIL, $\pi_{\text{MPPI}}$ is a sufficiently competitive generator in the adversarial dynamic which enables training the reward and value to completion.
4.4 Benchmarking – Q3
While we have shown theoretical justification for the classification of MPAIL as an AIL algorithm, we perform additional benchmarking experiments for empirical validation. We train GAIL, AIRL, and MPAIL on the navigation task and the cartpole task across varying quantities of expert demonstrations and random seeds as done in [12, 39]. While MPAIL uses an approximate prior model for the navigation task, we choose to learn a model during training of the cartpole task to demonstrate the generality of MPC and support future work on additional tasks. This is represented by the label, MPAIL (OM), indicating that there is a fully Online Model. Implementation details of the learned model can be found in Section D.3.
@@ 4.4 Benchmarking – Q3
Figure 6: Benchmarking Results. Top row rewards are computed across all demonstration quantities and seeds. Bottom row rewards are the average of the final 10 episodes computed across seeds.
@@
On the navigation task, MPAIL reaches optimality in less than half the number of interactions when compared to GAIL. We also observe MPAIL to train more stably than GAIL on this task. AIRL struggles to learn from the multimodal data (see Section 4.1). This navigation benchmark helps support MPAIL’s characterization as a model-based AIL algorithm as it is more sample-efficient when provided a prior model. On the cartpole task, expert demonstration data results in optimal-but-sparse state visitation and, equivalently, AIL reward signal. It is likely that, due to online dynamics model learning, MPAIL requires more exploratory interactions to combat a large local minima induced and reinforced by sparse discriminator reward, model bias, and task dynamics.
@@ 5 Conclusion
In this work, we formalize the Planning-from-Observation (PfO) setting, which, to the best of our knowledge, has not been previously named or addressed in an end-to-end manner. This is enabled by our introduction of one of the first complete algorithmic frameworks for PfO which unifies IRL and MPC, named MPAIL. We have shown that this formulation is theoretically and empirically justified and yields numerous benefits toward interpretable robot learning from observation. We find evidence that the integration of an online planner demonstrates substantial promise in robustness even when far outside of the data support of the expert. Compared to existing model-free AIL, MPAIL demonstrates empirical connections to model-based RL through improved sample-efficiency with a prior model. Towards real-world evaluation of AIL methods, we further deploy MPAIL through a Real-Sim-Real pipeline. From a single partially observable demonstration, MPAIL enables interpretable model-based planning even when learning from ambiguous expert data.
@@
MPAIL is derived from, and naturally admits, abstractions from Model Predictive Control, model-based RL, and imitation learning. Thus, its open-source implementation aims to reflect this and offers common ground for instantiating the many possible extensions to adjacent work through these connections: off-policy value estimation for improved sample efficiency [39], policy-like proposal distributions and latent dynamics for scaling MPPI [20], model-free and model-based reward blending for alleviating model bias [26], and much more. We envision that this work can provide a theoretically and empirically justified foundation for future work at the intersection of MPC, RL, and Imitation Learning.
6 Limitations
As noted in Section 3, the implications of the temperature decay remains to be investigated. We suspect that there may be a more adaptable approach towards scheduling or balancing its effects similar to well-studied entropy-regularized RL methods.
Vanilla MPPI is also known to struggle with high-dimensional tasks [40]. We direct the interested reader to Section D.6 to view results and discussion on the higher-dimensional ($\mathcal{S}\subseteq\mathbb{R}^{60}$; $\mathcal{A}\subseteq\mathbb{R}^{8}$) Ant task. However, vanilla MPPI still demonstrates promise towards scaling MPAIL through learned walking behavior. We believe that a biased sampling distribution, integration of existing MPPI optimization improvements [40], and latent state planning [20] offer promising solutions for scaling MPAIL to higher dimensions.
Consistent with recent findings on the limited reliability of simulation-based AIL when applied to human demonstrations [31], our experiments further highlight that real-world deployment of AIL would benefit from additional investigation [19]. In particular, choosing appropriate evaluation models can be cumbersome, especially in settings without clear performance metrics [41].
As AIL is fundamentally an interactive learning algorithm, MPAIL benefits from real-world RL approaches which seek to make training in the real-world practical. This would help alleviate losses in performance due to reward entanglement with environment dynamics by precluding the Real-Sim-Real alternative.
As observed in Footnote 2 and in our real-world experiments, un-trained regions of cost or reward are still capable of leading the agent astray. Albeit, this occurs at far lesser rates and impact than policy networks due to MPPI’s connections to optimal control and disturbance rejection. Further regularization or structural insights into training the discriminator are promising directions towards further robustifying these networks.

### v2

@@ 4 Experimental Results
Advancements in IRL and AIL continue to demonstrate improvements in sample efficiency. However, there remains an apparent gap with applications to robot learning. In addition to fundamental investigations about the MPAIL algorithm itself, our experiments target evaluations critical towards real-world robustness and generalization.
Without a policy, vanilla MPPI possesses no “memory” about actions taken in previous episodes, save for those implied through the value $V_{\phi}(s)$. And as discussed in Section 3, MPPI’s approximate online optimization is more practical for robustness and generalization but potentially less accurate for producing policies. These novelties raise a critical question about planning-based AIL; are policies solved online through planning sufficient as adversarially generative policies? We find that MPAIL indeed trains an effective imitator, provoking our follow-up questions:
@@
    What is the advantage of deploying an AIL planner over an AIL policy?
@@
    How does MPAIL help enable real-world planning capabilities from observation?
@@
Simulated Navigation Task
Figure 3: Four Expert Trajectories in Navigation Task. Cars initialized around $(0,0)$.
@@
For simulated evaluation, we design a navigation task with a 10-DoF vehicle. Reward is proportional to the negative squared distance to $(10,10)$. Initial poses are within 1 m of $(0,0)$. The state is 12-dimensions: position, orientation, linear velocity, and angular velocity. Actions include target velocity and steering angle. MPAIL plans using an approximate prior model, the Kinematic Bicycle Model (Han et al., 2024b). This approximate model is not tuned to the agent dynamics. For instance, slipping and suspension dynamics occur in simulation but are unmodeled.
@@
The expert for this environment is obtained through running PPO on a known reward (see Section D.2 for details) (Schulman et al., 2017b). At convergence, the optimal policy occasionally circles near the goal instead of remaining on it (see Figure 3). We choose to use the circling demonstrations as expert data, because it is a more challenging behavior to imitate. This is corroborated by (Orsini et al., 2021), who stress that demonstrator suboptimality and multimodality is a critical component in algorithm evaluation towards practical AIL from human data.
@@
This circling behavior acts as a critical “distractor mode”. In training, the policy may begin to only circle around the origin. If the AIL algorithm is not able to sufficiently explore, training collapses such that the policy continuously circles the origin, unable to return to the expert distribution which requires the circling behavior to occur around the goal. For instance, we find that AIRL is unable to successfully learn both behaviors likely due to the instability introduced by its logit shift (Orsini et al., 2021).
4.1 Out-of-Distribution Robustness Through Planning – Q1
@@
Figure 4: Comparison of policy-based and planning-based AIL in Out-of-Distribution (OOD) states. Agents trained on the navigation task (Section 4) are placed uniformly with random orientation between a 40 $\times$ 40 m box centered on $(0,0)$. The policy and planner are run for 100 timesteps in the environment. Data support of the expert exists mainly between $(0,0)$ and (10,10)(10,10) [^2] Note that the state space is 12-dimensional; expert data support is extremely sparse in this environment.. Quantitative evaluation of this experiment can be found in Figure 5. A comparison which includes a learned dynamics model can be found in Figure 8
@@
Figure 5: OOD Navigation Evaluation. Agent initial poses vary from In-distribution (ID) to OOD relative to the expert data and are plotted with their final reward after 100 timesteps. Metric from (Liu et al., 2020) (see Section D.2).
@@
When deploying learning-based methods to the real-world, reliable performance in out-of-distribution (OOD) states are of critical importance, especially in imitation learning when expert data can be extremely sparse. We show that planning-based AIL (MPAIL) improves generalization capabilities when OOD. In this experiment, we use the simulated navigation environment but expand the region of uniformly distributed initial positions and orientations from a $\times1$ square to a large $\times40$ m square around the origin (Figure 4). The policy-based approach is represented by GAIL, as AIRL does not meaningfully converge in the navigation task (Section 4). Only four expert trajectories are used in training.
We find that planning-based AIL generalizes to significantly more states than policy-based AIL when outside the support of expert data. In this experiment, the planner’s horizon is a maximum of 3 meters. As a result, the task horizon may be up to 15 times longer than the planning horizon. Evidently, a planner could not navigate to the goal if the learned optimization landscape (induced by cost $c_{\theta}$ and value $V_{\phi}$) did not also generalize to OOD states. These results suggest a fundamental limitation of current AIL approaches and their single policy solution. The trained reward and value are inefficiently underutilized in policy-based AIL and not utilized at all on deployment. By contrast, MPAIL re-introduces the reward and value online to solve for new policies each moment in time. These results illustrate that generalization in AIL is substantially improved through reward deployment in addition to reward learning.
4.2 Real-Sim-Real Navigation from a Single Observation – Q2
@@
Real-world evaluation of AIL is currently challenging. RL-like interaction efficiency renders training in simulation more practical than in the real-world (Tai et al., 2018), but demonstrations must realistically still be from the real-world. Nonetheless, it is imperative to evaluate AIL methods on real-world suboptimal data and hardware since results tend to diverge significantly from ideal settings, synthetic data, and simulation (Orsini et al., 2021; Tsurumine and Matsubara, 2022).
Our hardware experiment evaluates GAIL, IRL-MPC, and MPAIL through Real-to-Sim-to-Real: 1) a single partially observable (position and body-centric velocity) trajectory is collected from the real-world subject to sensor noise, 2) the method is trained using interactions from simulation, finally 3) the method is deployed zero-shot to the real-world for evaluation. This experiment uses a small-scale RC car platform with an NVIDIA Jetson Orin NX (Srinivasa et al., 2023). For IRL-MPC, the reward and value is trained through GAIL, which subsequently requires hand-tuning for deployment on MPPI (Triest et al., 2023).
We remark that the direction of travel cannot be uniquely determined by a single state $s$ due to the partially observable body-centric velocity. Only with the state-transition $(s_{E},s_{E}^{\prime})$ is it possible to deduce the direction of travel. This property is detailed in Section D.1. Partial observability and state-transitions play critical roles in the recovery of a cost function for this task, presenting a necessary challenge towards practical AIL (Orsini et al., 2021) and scalable Learning-from-Observation (LfO) (Torabi et al., 2019b).
Reward $r(s,s^{\prime})$ Policy Optimizer Deployment
GAIL $\log(D)$ PPO (Offline) Policy
AIRL $\log(\frac{D}{1-D})$ PPO (Offline) Policy
IRL-MPC $\log(D)$ PPO (Offline) Planner
MPAIL $\log(\frac{D}{1-D})$ MPPI (Online) Planner
Table 1: Summary of Baselines. The discriminator is denoted $D\coloneqq D(s,s^{\prime})$. Value estimation is performed via GAE-$\lambda$ (Schulman et al., 2018) for all methods.
We find that MPAIL is able to qualitatively reproduce the expert trajectory with an average Relative Cross-Track-Error (CTE (Rounsaville et al., )) of 0.17 m while traveling an average of 0.3 m/s slower. In addition, Figure 6 illustrates a key advantage of planning-based AIL. By granting access to the agent’s optimization landscape, MPAIL significantly improves on the interpretability of agents trained through observations of ambiguous and complex human data when compared to black-box policies. Note the lower costing of on-track trajectories and final plan.
GAIL does not reliably converge to the expert even in training. During deployment, GAIL’s policy consistently veers off-path or collapses into driving in a circle. Various starting configurations were attempted without success. While literature on the evaluation of AIL methods in the real-world are sparse, we find that AIL policies can be extremely poor performing in the real-world, as corroborated by (Sun et al., 2021). A more detailed discussion is provided in Section D.1.
IRL-MPC acts as a middle-ground between MPAIL and GAIL; the learned reward and value are exactly the same as GAIL’s and thus differs by the deployment of the reward through planning. IRL-MPC’s improvements over GAIL provides evidence that: (i) model-based planning can grant robustness to a model-free reward and, (ii) despite GAIL’s poor performance, the learned reward was still meaningfully discriminative and suggests a failure of the policy to arrive at a solution under the reward. On the other hand, IRL-MPC diverges from MPAIL by mainly learned reward and value. As a result, we find that online policy optimization through $\pi_{\text{MPPI}}$ induces a more competitive adversarial dynamic than offline policy optimization as in actor-critic RL. In this case, the end-to-end inclusion of the planner enables training the reward and value to completion.
@@
Table 2: Evaluation of Real Experiment. Relative Cross-Track Error (CTE) and speed are computed over the best five laps. AIRL is excluded as it does not exhibit meaningful behavior even in simulation.
Meanwhile, MPAIL’s success and IRL-MPC’s improvement over GAIL is attributed to model-based planning capabilities. If the robot should find itself away from the expert distribution, the online planner enables the agent to sample back onto the demonstration. In this sense, an MPC-based (or any online-optimizing) agent brings control-theoretic disturbance rejection online. Policies, on the other hand, are far more susceptible to erratic behavior in the real-world due to open-loop action prediction. This becomes especially important when the demonstration data is severely under-defined as in this partially observable setting, which results in ambiguous reward signal in most states in the environment; recall that low discriminator confidence is reflected by low magnitude logit $f_{\theta}$ (cost) predictions. Real experiment cost values are in the range of (0.022,0.0180)(-0.022,-0.0180) whereas cost values in benchmarking runs with synthetic demonstrations are in the range of (3,3)(-3,3). On real data, the discriminator also required more frequent updates to provide more reliable signals (see Table 3).
4.3 Efficiency – Q3
@@
Figure 7: Benchmarking Results. Top row rewards are computed across all demonstration quantities and seeds. Bottom row rewards are the average of the final 10 episodes computed across seeds. See Figure 11 for de-aggregated plots.
As mentioned at the beginning of Section 4, MPAIL does not possess a persistent “policy” as it employs a zeroth-order optimization with bootstrapped value estimates (i.e. infinite-horizon MPPI) to instead resolve these policies online. While we have shown that this “deconstructed policy” significantly improves generalization and robustness, concerns regarding interaction efficiency may arise due to the lack of gradient-based value optimization as in actor-critic policy optimization (Schulman et al., 2017b). Here, we perform additional benchmarking experiments to evaluate these hypotheses.
We train GAIL, AIRL, and MPAIL on the navigation task and the cartpole task across varying quantities of expert demonstrations and random seeds as done in (Ho and Ermon, 2016; Kostrikov et al., 2018). While MPAIL uses an approximate prior model for the navigation task, we choose to learn a model during training of the cartpole task to demonstrate the generality of MPC and support future work on additional tasks. This is represented by the label, MPAIL (OM), indicating that there is a fully Online Model. Implementation details of the learned model can be found in Section D.3.
@@
On the navigation task, MPAIL reaches optimality in less than half the number of interactions when compared to GAIL. We also observe MPAIL to train more stably than GAIL on this task. AIRL struggles to learn from the multimodal data (see Section 4). In addition to alleviating concerns regarding efficiency, these results support MPAIL’s characterization as a model-based algorithm as it is more sample-efficient when provided an (approximate) prior model.
On the cartpole task, expert demonstration data results in optimal-but-sparse state visitation and, equivalently, AIL reward signal. It is likely that, due to online dynamics model learning, MPAIL (OM) requires more exploratory interactions to combat a large local minima induced and reinforced by sparse discriminator reward, model bias, and task dynamics. Similar performance between MPAIL and AIRL may also suggest that GAIL benefits from its inherent reward bias on this task (Kostrikov et al., 2018). Importantly, MPAIL attains comparable performance while maintaining the benefits of model-based planning, such as interpretability, transferability, and robustness.
Additional Results
In the Appendix, additional experiments are performed to holistically evaluate MPAIL for the purposes of real-world deployment and robot learning. Wall clock time comparisons, architecture, hyperparameters, model ablations, proofs, de-aggregated benchmark results, and more experiment details and discussion can be found in the appendix. We briefly highlight some key auxiliary results:
Wall Clock Time. Our timing evaluations reveal that “inference” and training times of MPAIL can be faster or slower than GAIL (PPO) depending mostly on the number of MPPI iterations per step and the planning horizon $H$. For $H=10$, MPAIL is about $10\%$ faster than GAIL at 2 iterations. At 5 iterations, the same horizon is about $10\%$ slower. More settings are evaluated and shown in Figure 10.
MPAIL (OM) — OOD Experiment. Does OOD robustness hold when the dynamics model is also learned? Figure 8 compares the final states of the OOD evaluation in Section 4.1 with MPAIL (OM) included. Note that MPAIL (OM) has access to the same amount of information as the policy-based method. Nevertheless, there is clear improvement in OOD performance by MPAIL (OM) when compared to policy networks. These results crucially suggest that generalizability advantages of the deconstructed policy extends to learned dynamics models in addition to reward and value.
Figure 8: Out-of-Distribution (OOD) performance evaluation (as in Figure 4) at $t=100$ with MPAIL (OM) and Behavior Cloning (BC) included. **BC requires access to expert actions and is not an LfO baseline. MPAIL (OM) indicates that a dynamics model is learned online and used for planning in MPAIL. Without a prior model, MPAIL (OM) is limited to the same amount of total information as GAIL; any improvement in OOD generalization over a policy network is purely a result of learning a deconstructed policy for online planning. Details of MPAIL (OM) in Section 4.3 and Section D.3.
Even in the case of Behavior Cloning (BC), which is directly supervised with access to expert actions, generalization of the policy network appears random and unpredictable. For policy networks (BC and GAIL), there are numerous agents that are initialized near the expert distribution (within red highlight) but remain there. These agents were often merely initialized facing away from the goal, demonstrating that policy networks tend to learn incredibly brittle representations. Most MPAIL (OM) agents that are seen far away from the goal will continue to slowly arrive at the goal, visually evident by their orientations directed towards the goal. Contrasting with MPAIL, MPAIL (OM) agents tend to follow longer, less optimal paths when strongly OOD. This is likely a result of the learned dynamics being OOD and disrupting planning.
@@ 5 Conclusion
This work adopts an imitation learning setting familiar to humans and animals in which: (1) expert actions are not known, (2) few demonstrations are observed, and (3) the agent infers intent and improves through interaction. While work in Inverse Reinforcement Learning (IRL) and Adversarial Imitation Learning (AIL) from Observation continue to advance these goals, their applications to real-world robots lag behind. To bridge this gap, we observe that connections to model-based planning offers potential towards (a) improved efficiency and transfer, (b) safe and steerable design, and (c) robustness through online optimization. We call this problem setting Planning-from-Observation (PfO). We address PfO by introducing planning-based AIL as a unification of IRL and MPC. Model Predictive Adversarial Imitation Learning (MPAIL) is then introduced as an implementation of planning-based AIL, where a planner is continually improved through cost and value learning.
Towards robot learning applications, we conducted three elucidating evaluations. In Section 4.1, we reveal that reward deployment—not only reward learning as in current AIL—is critical towards generalizable imitation learning. Re-introducing the learned reward online alleviates the burden on policies to generalize and requires the reward, or intent, to generalize instead. In Section 4.2, we see that reward deployment must be met with online optimization for real-world robustness. Especially in partially observable AIL, it may not be sufficient to utilize a planner only during deployment, but the planner should also be included in the learning process to train the reward to completion. From a single partially observable demonstration in real-world navigation, we find that MPAIL is the only successful imitator when compared to policy-based AIL and IRL-MPC. MPAIL also employs representations (e.g. model, reward, value) which grant direct access to the agent’s optimization landscape and thus decision-making process—a powerful prerequisite towards safe and interpretable robot learning. Finally, in Section 4.3, interaction efficiency benchmarks favorable to MPAIL with an approximate prior model address concerns regarding MPPI’s zeroth-order policy optimization and empirically supports MPAIL’s characterization as a model-based algorithm.
@@
MPAIL is derived from, and naturally admits, abstractions from Model Predictive Control, model-based RL, and imitation learning. Thus, its open-source implementation aims to reflect this and offers common ground for instantiating the many possible extensions to adjacent work through these connections: off-policy value estimation for improved sample efficiency or offline learning (Kostrikov et al., 2018), policy-like proposal distributions and latent dynamics for scaling MPPI to higher dimensional spaces (Hansen et al., 2024; Zhu et al., 2025), model-free and model-based reward blending for alleviating model bias (Bhardwaj et al., 2021), diffusion-inspired MPPI for improved online optimization (Xue et al., 2024), and much more (Han et al., 2026). We envision that this work can provide a theoretically and empirically justified foundation for future work at the intersection of MPC, RL, and Imitation Learning.
Reproducibility
Included in the supplementary material is the code for MPAIL, developed with the intention to be released as a standalone package. Configuration files and hyperparameters for experiments are almost entirely “flat” (viewable in the code) and are as presented in Table 3. Simulation environment is built with Isaac Lab (Mittal et al., 2023) and one of its extensions, Wheeled Lab (Han et al., 2025). Instructions for installation can be found in their respective GitHub repositories. Hardware deployment code is also included for the (open-source) platform, MuSHR (Srinivasa et al., 2023); an example script detailing how to save and load an MPAIL planner onto the hardware can be found in the supplementary as well. More details explaning source code organization, as well as experiment videos, can be found in the README.md. We welcome efforts towards reproducibility and encourage practitioners to correspond regarding difficulties.